OpenStudy (anonymous):

The Ruffs are planning to buy an above-ground swimming pool shaped as a regular octagon. The radius of the octagon is 9 feet. To the nearest tenth, find the area of the surface of the water in the pool. Possible answers: A) 458.2 ft2 B) 553.1 ft2 C) 94.8 ft2 D) 229.1 ft2 Thanks!

OpenStudy (anonymous):

A=r^2*2.828

OpenStudy (anonymous):

So A=(9)^2*2.828=229.068

OpenStudy (dumbcow):

octagon can be split into 8 equal isosceles triangles Area of octagon will be 8X area of each triangle 360/8 = 45 top angle opposite of base is 45 degrees 180(8-2) = 1080, this is sum of interior angles 1080/8 = 135 135/2 = 67.5 bottom 2 angles opposite of radii are 67.5 degrees to find area of triangle we need base length(x) and height(h) use trig ratios h = 9sin(67.5) x = 18cos(67.5) Area = 8*(1/2)*(9sin(67.5))*(18cos(67.5)) Area = 648*sin(67.5)*cos(67.5) Area = 229.1